OpenStudy (anonymous):

Write the polynomial in simplest form with roots 4, 5i, and -5i.

OpenStudy (aum):

Take the complex roots first. You just need one root because complex roots always occur in pairs. x = 5i square both sides: x^2 = (5i)^2 = -25 (x^2 + 25) = 0 will have the roots 5i and -5i Now multiply (x^2 + 25) by (x-4) to get the polynomial.

OpenStudy (anonymous):

since the roots are 4, 5i, and -5i. the polynomial is of the form (x-4)(x+5i)(x-5i)

OpenStudy (anonymous):

Thank you!

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

Could you possilbly do another example? I dont really understand it

OpenStudy (aum):

Well, another way to do the same problem is what matricked suggested above. If three roots, a, b and c are given, then the simplest polynomial with those three roots is: (x-a)(x-b)(x-c). So here it is (x+5i)(x-5i)(x-4) First multiply (x+5i)(x-5i). Use the identity (a+b)(a-b) = a^2 - b^2 So (x+5i)(x-5i) = x^2 - (5i)^2 = x^2 + 25 Then multiply (x^2+25)(x-4)

OpenStudy (anonymous):

Like foils, thank you! is this for all polynomial things, no matter what roots are given?

OpenStudy (aum):

Yes. If they given you any number of roots and ask for the simplest polynomial with those roots, then just proceed as shown. This works because if 'a' is a root it implies (x-a) is a factor.

OpenStudy (aum):

So if a, b, c, d, e are roots then (x-a), (x-b), (x-c), (x-d), (x-e) are factors and you can multiply them all and find the polynomial. Example find the simplest polynomial with roots 2 and -3. Answer: (x-2)(x- -3) or (x-2)(x+3). FOIL it and simplify to get x^2 + x - 6.